OpenStudy (anonymous):

how do I solve for t in the equation x=9ln(3t) ?

OpenStudy (anonymous):

x/9 = ln(3t) 3t = e^(x/9) t = 1/3*e^(x/9)

OpenStudy (anonymous):

To solve for t you raise e to the power of each side.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[e^x = e^{9ln(3t)}\]\[\implies e^x = (e^{ln(3t)})^9\]\[\implies (3t)^9 = e^x\]\[\implies 3t = e^{\frac{1}{9}x}\]\[\implies t = \frac{1}{3}e^{\frac{1}{9}x}\]